Some systems are experiencing issues

About This Site

OFFICE HOURS: Office Hours will be held virtually until further notice. Please see https://www.rc.fas.harvard.edu/training/office-hours/

GETTING HELP

https://docs.rc.fas.harvard.edu | https://portal.rc.fas.harvard.edu/rcrt/submit_ticket | Email: rchelp@rc.fas.harvard.edu


Status page for the Harvard FAS Research Computing cluster and other resources.

Please scroll down to see details on any Incidents or maintenance notices.

Maintenance
Home Directories unavailable All Day July 20th, 2020 - Affects most services

**Home directory physical server move – All day event **

Home directories will be physically moved. This will require the stoppage of all jobs and cluster logins as accounts cannot operate without a home directory.

We are targeting a single day (~ 5am-8pm) on July 20th.

High Impact: All cluster jobs, cluster login, VDI/OOD login, home directory mounts, software modules, Portal, and new account signups/approvals

Does not directly affect SPINAL, MiniLims, Windows-only environments (except if home directories have been manually mounted) but some impact is possible.

Please read details and any updates here: https://www.rc.fas.harvard.edu/isilonmove-2020

Past Incidents

13th July 2020

Primary Lab Storage - Boston rcnfs12 3 disk failure

rcnfs12 suffered a 3 disk failure on a RAID 6 set. We are working to reconstruct. This impacts the following shares:

ccfa kuang_lab hsphfs1_breakbulk2 samuel_lab philipkim_lab desmarais_lab scharf_lab piquet_lab hopkins_lab

Samba Cluster samba cluster partial outage

One of the samba hosts is in a bad state. We are working to resolve. samba connections to RC storage may hang until we fix.

11th July 2020

BosLFS boslfs stuck OST's

boslfs has several stuck OST's. We are power cycling.

HolyLFS holylfs stuck OST's

Several OST's are stuck on holylfs, we are cycling the system.

9th July 2020

Holystore01 holystore01

Due to improper user jobs, holystore01 is wedged. We are investigating.

  • holystore01 should be back

  • 8th July 2020

    HolyLFS holylfs

    holylfs is experiencing slow performance. We are investigating the root cause.

  • holylfs is back

  • Holystore01 holystore01 latency

    OST's on holystore01 are stuck causing latency issues. We are cycling to fix.

    7th July 2020

    Portal Password reset not working

    The Portal's password reset function is not working. We are investigating. Updates to follow.

  • The issue has been resolved.

  • 4th July 2020

    HolyLFS holylfs stuck OST's

    Several OST's on holylfs are stuck. We are cycling.

    For issues not shown here, please contact FASRC via
    https://portal.rc.fas.harvard.edu or email rchelp@rc.fas.harvard.edu